The augmented matrix
1 3 2 4
2 -1 5 -2
-3 1 6 1
Represent the systems of equations.

Answer by checkley71(8403)

X+3Y+2Z=4
2X-Y+5Z=-2
-3X+Y+6Z=1
THESE ARE THE 3 EQUATIONS IN THE MATRIX.